# Exploring the generalizability of the optimal 0.234 acceptance rate in random-walk Metropolis and parallel tempering algorithms

### v3

@@ 3.3.1 The “Rough Carpet” Target Density
where the single component density $f_{3}$ is a one-dimensional density with three modes at $(\mu_{1},\mu_{2},\mu_{3})=(-5,0,5)$ and $\mathcal{N}(x|\mu_{i},1)$ indicates the density of the univariate Gaussian distribution at point $x$ with mean $\mu_{i}$ and variance 1. This forms a “rough carpet”-like appearance for the target distribution when we take the product of this single-component density over many dimensions (see the visualization of the first two components in Figure 3 for an example). Since we have $3^{d}$ modes, the density values at each mode may not be very high. Yet, since this is a i.i.d. product form, the theoretical result of 0.234 being the optimal acceptance rate should also apply.
@@ 3.3.2 The “Rough Carpet” Target Density: Inhomogeneously Scaled Components
Note that $f_{5}=f_{4}$, so this density $\pi_{5}$ is a more general form of $\pi_{4}$ by including inhomogeneous component-wise scaling factors $C_{i}$, where $C_{i}>0$ and $\mathbb{E}(C_{i})=1$ and $\text{Var}(C_{i})<\infty$. This inhomogeneously-scaled i.i.d. product target density was discussed by [33], which proved the extension of the 0.234 asymptotically optimal acceptance rate to this case. We independently sample the component scaling factors $C_{i}\sim\text{Uniform}[0.02,1.98]$.
@@ 3.4 Rosenbrock Target Densities
The multimodal experiments in Section 3.3 demonstrated the effectiveness of the 0.234 acceptance rate for multimodal distributions, where the main challenge is mode isolation. In practice, however, many posterior landscapes are not only multi-peaked but also have nonlinear correlation structure, providing challenges for local proposals even when there is only a single global mode. We examine this setting by sampling three different extensions of the two-dimensional Rosenbrock target density to $n$-dimensions [11, 9, 27], each of which have different curvature and therefore different difficulty from a MCMC perspective.
@@ 3.4.1 Full Rosenbrock
The original 2-D Rosenbrock resembles a long banana-shaped valley whose floor follows the parabola $x_{2}=x_{1}^{2}$. As stated in Pagani et al. [27], as $d$ increases, the long, narrow ridge that makes sampling challenging becomes more concentrated around the mode, decreasing the problem’s difficulty. However, the variance of $x_{d}$ increases steeply as each new random variable is directly dependent on the squared value of the previous variable.
@@ 3.4.2 Even Rosenbrock
The Even Rosenbrock partitions the state into $\tfrac{d}{2}$ independent 2-D Rosenbrock blocks. Because the blocks do not interact, the global target is the product of smaller bananas; the joint density still exhibits the same curved valleys within each block but lacks the long-range, serial dependence of the Full Rosenbrock. Consequently, the optimal proposal scale can adapt locally to each block, and we expect higher ESJD (and hence slightly lower optimal acceptance rates) than in the fully coupled case.
@@ 3.4.3 Hybrid Rosenbrock
The $n$-dimensional Hybrid Rosenbrock distribution introduced by Pagani et al. [27] has the following kernel for $\mathbf{x}\in\mathbb{R}^{d}$, where $\mu,x_{j,i}\in\mathbb{R};a,b_{j,i}\in\mathbb{R}^{+}(\forall j,i)$:
@@ 3.4.4 Simulations and Results
For our Rosenbrock simulations, we found that our initial settings of 200,000 Markov chain steps and 20 seeds were insufficient to produce meaningful trends, so we ran more simulations for more steps with the Rosenbrock targets. In this section, we continue using a Gaussian proposal with 40 different variance values. However, we run each simulation for 1,000,000 steps and we repeated each experiment with 30 different seeds instead of the usual 20. For the Full Rosenbrock, we select $d\in\{5,10,20,30\}$. For the Even Rosenbrock, we select $d\in\{4,10,20,30\}$. For the Hybrid Rosenbrock, we select $(n_{1},n_{2})\in\{(3,2),(5,2),(7,3),(8,4)\}$ which gives total dimension $d\in\{5,9,19,29\}$ respectively. We also provide a plot of the ESJD versus acceptance rate trend for the 2-dimensional Rosenbrock. All other plots of ESJD versus acceptance rate trends for these distributions are in Appendix A Figures 16, 17, and 18 for plots of the trend of ESJD with acceptance rate.
@@
Our experiments with the Rosenbrock family of targets, presented in Table 4, demonstrate that the 0.234 heuristic remains effective even for distributions with non-linear dependencies and curvature. For the Full Rosenbrock ($\pi_{8}$) and Hybrid Rosenbrock ($\pi_{10}$) targets, the optimal acceptance rate robustly converges to approximately 0.234 as the dimension increases. Although these targets are not i.i.d. and possess a highly anisotropic banana-shaped geometry, they should satisfy the general criteria from Sherlock [38] (stated as Assumption 2.2) in the high-dimensional limit. As $d$ grows, the complex local correlations become less influential relative to the global structure, allowing the target’s geometric properties to stabilize in the manner required by Assumption 2.2. On the other hand, the Even Rosenbrock ($\pi_{9}$) has a significantly lower optimal acceptance rate that increases slowly with dimension. This seems reasonable, as $\pi_{9}$ is a product of independent 2-dimensional Rosenbrock blocks. The proposal scale can then be fit to the local geometry of each 2-dimensional block where the high-dimensional scaling theory is less applicable.
@@
The result for the 2-dimensional Rosenbrock target, shown in Figure 5, would add support for this hypothesis. Unlike the higher-dimensional cases, the ESJD does not exhibit a clear peak around a non-zero acceptance rate. Instead, it appears to increase monotonically as the acceptance rate approaches zero. This behaviour indicates that for this low-dimensional, highly anisotropic target, the highest efficiency is achieved by proposing extremely large jumps. Such jumps are almost always rejected, but the rare accepted jump is so large that it dominates the ESJD calculation.
@@ 3.5 Non-smooth Target Densities: Continuous Hypercube
Now, consider a RWM using a Gaussian proposal and a target density still of the product form $\pi(\mathbf{x})=\prod_{i=1}^{d}f(x_{i})$, but where the individual component density $f$ is Uniform$[a,b]$.
@@
As the dimension $d$ increases, the algorithm’s performance converges to the theoretical 0.1353 value. In this specific case, the shrinking optimal acceptance rate might be because of the unique zero or non-zero form of the target density beyond the hypercube boundary. As discussed by [25], whether a proposed move is accepted depends completely on whether the proposed state is inside the hypercube. Subsequently, when using a Gaussian proposal, the probability that at least one proposed component will lie outside the boundary increases, leading to a higher probability of instant rejection.
@@ 3.6 Neal’s Funnel Target Density
For our experiments, we follow the standard parameterization from Neal [22]. We test this target in dimensions $d\in\{5,10,20,30\}$. Due to the difficulty of sampling this target, we increased our simulation budget significantly, using $1{,}000{,}000$ iterations per simulation (across 40 seeds, 40 million total Metropolis steps), a burn-in of $50{,}000$ steps, and averaging the results over $30$ random seeds for each data point. We continue using the Gaussian proposal with variance as described in (8).
@@ 4.1 Parallel Tempering Method
The parallel tempering method [10], also known as the Metropolis-coupled Markov chain Monte Carlo algorithm or the replica exchange method, is a version of the Metropolis algorithm that specialises in sampling from multimodal distributions. In this case, we sample from a target distribution with density $\pi(\mathbf{x})$ on a state space $\mathcal{X}$. We define a sequence of tempered target densities $\pi^{\beta_{j}}(\mathbf{x})$ where $10\leq\beta_{n}<\beta_{n-1}<\dots<\beta_{1}<\beta_{0}=$ 1 are the inverse temperature values. We require that $\pi^{\beta_{0}}(\mathbf{x})=\pi(\mathbf{x})$ (that is, that $\beta_{0}=1$). Typically, the tempered target densities are simply powers of the original density, and we assume this for all parallel tempering experiments. Then, we run a Markov chain for each of the $n+1$ values of $\beta$, each with its own tempered target density. Eventually, the $\beta_{0}=1$ chain, also called the “cold” chain, with its stationary density $\pi^{\beta_{0}}(\mathbf{x}_{0})$ should correspond to the original target density of interest $\pi(\mathbf{x})$, and the idea is that using information from other chains with higher temperatures (lower $\beta$) will speed up the mixing in the cold chain.
@@
4.1.1 Optimal Scaling for Parallel Tempering
@@
    The target density is an i.i.d. product $\pi(\mathbf{x})=\prod_{i=1}^{d}f(x_{i})$ where $f$ is a single-dimensional component density.
@@ 5.3 Results
The empirical optimal swap acceptance rates for the three target distributions are summarized in Table 6, and sample ESJD curves for $d=30$ shown in Figure 8. The ESJD curves for PT are visibly more jagged than their RWM counterparts. This might be because the estimation of ESJD in this context is more susceptible to Monte Carlo error even with a large number of simulation steps, since the efficiency or acceptance rate of a swap move depends on the joint stationary distribution of two adjacent chains, and this is repeated for each temperature construction. Nonetheless, clear trends emerge from the results.
@@
Notably, the Rough Carpet target ($\pi_{4}$) shows a significant deviation. The optimal swap acceptance rate is consistently higher than 0.234 and increases away from the theoretical value with increasing dimension. This is surprising because the Rough Carpet target does satisfy the i.i.d. product form and other conditions of Assumption 4.1 by Atchadé et al. [1]. The failure of the 0.234 rule here could be because the extreme multimodality of the target could reduce the practical utility of the ”hot” chains. The target possesses $3^{d}$ modes, and its probability mass is spread thinly across this vast landscape. For a hot chain where the inverse temperature $\beta$ is close to 0, the tempered density becomes extremely flat and loses the slight peak that a unimodal target would have. Consequently, it fails to efficiently sample from the specific modal regions that are of interest to the colder chains; the information flow from hot to cold chains is broken. To swap successfully more often, the temperature steps must be made smaller, which forces the swap acceptance rate to be higher.
@@ 6 Discussion
We have presented an extensive empirical investigation into the generalizability of the 0.234 acceptance rate heuristic for both Random-Walk Metropolis (RWM) and Parallel Tempering (PT) algorithms, using a variety of target and proposal distributions. Our results demonstrate that the 0.234 rule’s utility extends beyond the classical i.i.d. product setting of Roberts et al. [29]. The heuristic remains approximately optimal for: (i) non-Gaussian proposals that satisfy a spherical symmetry (shell) condition, such as the Laplace and Uniform distributions; (ii) multimodal targets such as the “rough carpet” i.i.d product with $3^{d}$ modes, and non-product targets constructed as a sum of isotropic components, such as our Three-Mixture model; and (iii) even for highly correlated, non-linear targets like the Full and Hybrid Rosenbrock in sufficiently high dimensions ($d\geq 20$). These findings provide empirical validation for the generalized theory of Sherlock [38], which predicts that as long as the target’s geometry is sufficiently regular and isotropic in the high-dimensional limit, the 0.234 rule should hold.
@@
That being said, there are cases where 0.234 does not seem as relevant. The main case is in very low dimensions; in all our experiments with varying target densities, the optimal acceptance rate strays from 0.234 below a certain threshold which varies for each target density. In particular, targets composed of independent, low-dimensional anisotropic blocks (Even Rosenbrock) or those that are globally anisotropic in low dimensions (2D Rosenbrock), have an optimal acceptance rate significantly below 0.234. In these cases, we hypothesize the high-dimensional averaging effect that stabilizes the target’s geometry has not occurred, and the sampler’s performance is dictated by the challenging local geometry. As the dimension increases, the law of large numbers mitigates the effect of any single scaling factor or local correlation, so the target becomes effectively less eccentric, complying with the theory of Sherlock [38], and the optimal acceptance rate recovers towards 0.234. On the other hand, the continuous hypercube and Neal’s funnel completely fail the 0.234 heuristic. Although the hypercube is continuously differentiable everywhere except on its boundary, which has zero measure, an intuitive argument is that as $d$ increases, the probability that the proposal has all the values inside decreases. As for the funnel, no single proposal scaling can be globally efficient, which is just a limitation of the RWM algorithm.
@@
Next, we discuss the optimal swap acceptance rate figure for parallel tempering and compare it with the optimal acceptance rate for random-walk Metropolis. On one hand, the results for the multivariate Gaussian and three-mixture density examples show that the 0.234 swap acceptance rate figure may be optimal even in lower dimensions. Although costly, if the practitioner is willing and able to construct the inverse temperature ladder iteratively with a spacing given by a swap acceptance probability, the spacing dictated by a 0.234 swap acceptance probability may be optimal for the practitioner. However, the theoretical 0.234 figure may not always be applicable: despite satisfying the i.i.d. conditions of the PT optimal scaling theory [1], the extreme multimodality ($3^{d}$ modes) of the “rough carpet” product density appears to render the hot chains ineffective. We hypothesize that because the probability mass is spread so thinly, the hot chains fail to sample from meaningful modal regions, breaking the flow of information to the cold chain. This forces the algorithm to use smaller temperature steps (and thus higher swap acceptance rates) to achieve mixing, representing a practical failure of the 0.234 PT rule.

### v4

@@ 3.3.1 The “Rough Carpet” Target Density
where the single component density $f_{4}$ is a one-dimensional density with three modes at $(\mu_{1},\mu_{2},\mu_{3})=(-5,0,5)$ and $\mathcal{N}(x|\mu_{i},1)$ indicates the density of the univariate Gaussian distribution at point $x$ with mean $\mu_{i}$ and variance 1. This forms a “rough carpet”-like appearance for the target distribution when we take the product of this single-component density over many dimensions (see the visualization of the first two components in Figure 3 for an example). Since we have $3^{d}$ modes, the density values at each mode may not be very high. Yet, since this is a i.i.d. product form, the theoretical result of 0.234 being the optimal acceptance rate should also apply.
@@ 3.3.2 The “Rough Carpet” Target Density: Inhomogeneously Scaled Components
Note that $f_{5}=f_{4}$, so this density $\pi_{5}$ is a more general form of $\pi_{4}$ by including inhomogeneous component-wise scaling factors $C_{i}$, where $C_{i}>0$ and $\mathbb{E}(C_{i})=1$ and $\text{Var}(C_{i})<\infty$. This is a special case of a class of inhomogeneously-scaled i.i.d. product target densities discussed by [34], which proved the extension of the 0.234 asymptotically optimal acceptance rate to this class. We independently sample the component scaling factors $C_{i}\sim\text{Uniform}[0.02,1.98]$.
@@ 3.4 Rosenbrock Target Densities
The multimodal experiments in Section 3.3 demonstrated the effectiveness of the 0.234 acceptance rate for multimodal distributions, where the main challenge is mode isolation. In practice, however, many posterior landscapes are not only multi-peaked but also have nonlinear correlation structure, providing challenges for local proposals even when there is only a single global mode. We examine this setting by sampling three different extensions of the two-dimensional Rosenbrock target density to $d$-dimensions [12, 10, 28], each of which have different curvature and therefore different difficulty from a MCMC perspective.
@@ 3.4.1 Full Rosenbrock
The original 2-dimensional Rosenbrock resembles a long banana-shaped valley whose floor follows the parabola $x_{2}=x_{1}^{2}$. As stated in Pagani et al. [28], as $d$ increases, the long, narrow ridge that makes sampling challenging becomes more concentrated around the mode, decreasing the problem’s difficulty. However, the variance of $x_{d}$ increases steeply as each new random variable is directly dependent on the squared value of the previous variable.
@@ 3.4.2 Even Rosenbrock
The Even Rosenbrock partitions the state into $\tfrac{d}{2}$ independent 2-dimensional Rosenbrock blocks. Because the blocks do not interact, the global target is the product of smaller bananas; the joint density still exhibits the same curved valleys within each block but lacks the long-range, serial dependence of the Full Rosenbrock. Consequently, we expect higher ESJD (and hence slightly lower optimal acceptance rates) than in the fully coupled case.
Our experiments only use a single scaling factor applied to the whole covariance matrix of the proposal, but it could be useful for a practitioner to note that it is possible to locally tune the proposal for each 2-dimensional Rosenbrock block to make the RWM algorithm more efficient on this distribution. For example, when using a Gaussian proposal, the covariance matrix could be set as a block diagonal matrix where each 2x2 block on the diagonal of the Gaussian covariance matrix is tailored to and achieves optimality on its own 2-dimensional Rosenbrock.
@@ 3.4.3 Hybrid Rosenbrock
The $d$-dimensional Hybrid Rosenbrock distribution introduced by Pagani et al. [28] has the following kernel for $\mathbf{x}\in\mathbb{R}^{d}$, where $\mu,x_{j,i}\in\mathbb{R};a,b_{j,i}\in\mathbb{R}^{+}(\forall j,i)$:
@@ 3.4.4 Simulations and Results
For our Rosenbrock simulations, we found that our initial settings of 200,000 Markov chain steps and 20 seeds were insufficient to produce meaningful trends, so we ran more simulations for more steps with the Rosenbrock targets. In this section, we continue using a Gaussian proposal with 40 different variance values. However, we run each simulation for 1,000,000 steps and we repeated each experiment with 30 different seeds instead of the usual 20. For the Full Rosenbrock, we select $d\in\{5,10,20,30\}$. For the Even Rosenbrock, we select $d\in\{4,10,20,30\}$. For the Hybrid Rosenbrock, we select $(n_{1},n_{2})\in\{(3,2),(5,2),(7,3),(8,4)\}$ which gives total dimension $d\in\{5,9,19,29\}$ respectively.
We also provide a plot of the ESJD versus acceptance rate trend for the 2-dimensional Rosenbrock. Note that for $d=2$, each of the three Rosenbrocks reduce to the standard Rosenbrock 2-dimensional density, assuming $\mu_{1}=1$ for the Even Rosenbrock and $n_{1}=2,n_{2}=1,a=\frac{1}{20},b=\frac{100}{20},\mu=1$ for the Hybrid Rosenbrock. All other plots of ESJD versus acceptance rate trends for these distributions are in Appendix A Figures 16, 17, and 18 for plots of the trend of ESJD with acceptance rate.
@@
Our experiments with the Rosenbrock family of targets, presented in Table 4, demonstrate that the 0.234 heuristic remains effective even for distributions with non-linear dependencies and curvature. For the Full Rosenbrock ($\pi_{8}$) and Hybrid Rosenbrock ($\pi_{10}$) targets, the optimal acceptance rate robustly converges to approximately 0.234 as the dimension increases. Although these targets are not i.i.d. and possess a highly anisotropic banana-shaped geometry, they should satisfy the general criteria from Sherlock [39] (stated as Assumption 2.2) in the high-dimensional limit. As $d$ grows, the complex local correlations become less influential relative to the global structure, allowing the target’s geometric properties to stabilize in the manner required by Assumption 2.2. On the other hand, the Even Rosenbrock ($\pi_{9}$) has a significantly lower optimal acceptance rate in low dimensions than the other Rosenbrocks, and the optimal acceptance rate increases slowly towards 0.234 with increasing dimension. This seems reasonable and the analysis of Beskos et al. [6] is useful here.
The Even Rosenbrock is a product of independent 2-dimensional Rosenbrock blocks, and for a small number of blocks, the standard high-dimensional scaling theory is less applicable. We elaborate on this in the next paragraph about the 2-dimensional Rosenbrock, but the key relevant message from Beskos et al. [6] is that in low dimensions, large proposed jumps and an acceptance rate closer to zero are most efficient, explaining the low optimal acceptance rate. As the dimension $d$ and thus the number of independent blocks grow, the dynamics transition from a Markov jump process towards a limiting diffusion. While the theory for such targets suggests a position-dependent diffusion speed and thus a “local” 0.234 rule [6, see point (iv), p. 2968], the growing number of independent blocks and therefore the growing number of “large” components may induce a law of large numbers effect where position-dependent variations in diffusion speed average out across the many blocks. This would result in an approximately constant global diffusion speed and therefore a global 0.234 acceptance rate being approximately optimal in high dimensions for the Even Rosenbrock.
@@
The result for the 2-dimensional Rosenbrock target, shown in Figure 5, provides empirical evidence for the low-dimensional sampling phenomena analyzed by Beskos et al. [6]. Unlike the higher-dimensional cases, the ESJD does not exhibit a clear peak around a non-zero acceptance rate. Instead, it appears to increase monotonically as the acceptance rate approaches zero.
As theoretically established by Beskos et al. [6], the optimal sampling strategy for this low-dimensional ridged target is not a diffusion process. They state that efficiency is maximized by proposing extremely large jumps using a large proposal scaling $\sigma$ and allowing the acceptance rate to go to zero. A heuristic explanation for this is that in this case, the squared distance of a proposal increases proportionally to $\sigma^{2}$ but the fraction of accepted proposals decreases proportionally to $\frac{1}{\sigma}$, and therefore ESJD increases proportionally to $\sigma$ until the proposal scale becomes comparable to the characteristic scale of the target distribution itself, at which point the acceptance probability begins to decay more rapidly. Taking the limit, as the thickness (scale $\varepsilon$ in Beskos et al. [6]) of the support of the target distribution goes to zero, the algorithm’s dynamics converge to a continuous-time Markov jump process rather than a diffusion. Our observation for the 2-dimensional Rosenbrock target, and by extension the low-dimensional Even Rosenbrock target (e.g., for $d=4$, which consists of two independent 2-dimensional blocks), is an instance of this theoretically-predicted behaviour.
@@ 3.5 Non-smooth Target Densities: Continuous Hypercube
[…]
@@
As the dimension $d$ increases, the algorithm’s performance converges to the theoretical 0.1353 value. In this specific case, the shrinking optimal acceptance rate might be because of the unique zero or non-zero form of the target density beyond the hypercube boundary. As discussed by [26], whether a proposed move is accepted depends completely on whether the proposed state is inside the hypercube. Subsequently, when using a Gaussian proposal, the probability that at least one proposed component will lie outside the boundary increases as dimension increases, leading to a higher probability of instant rejection when the proposal scaling factor independent of dimension is kept constant.
@@ 3.6 Neal’s Funnel Target Density
For our experiments, we follow the standard parameterization from Neal [23]. We test this target in dimensions $d\in\{5,10,20,30\}$. Due to the difficulty of sampling this target, we increased our simulation budget significantly, using 1,000,0001{,}000{,}0001 , 000 , 000 iterations per simulation (across 40 seeds, 40 million total Metropolis steps), a burn-in of 50,00050{,}00050 , 000 steps, and averaging the results over 303030 random seeds for each data point. We continue using the Gaussian proposal with variance as described in (8).
@@ 4.1 Parallel Tempering Method
The parallel tempering method [11], also known as the Metropolis-coupled Markov chain Monte Carlo algorithm or the replica exchange method, is a version of the Metropolis algorithm that specialises in sampling from multimodal distributions. In this case, we sample from a target distribution with density $\pi(\mathbf{x})$ on a state space $\mathcal{X}$. We define a sequence of tempered target densities $\pi^{\beta_{j}}(\mathbf{x})$ where $0\leq\beta_{n}<\beta_{n-1}<\dots<\beta_{1}<\beta_{0}=1$ are the inverse temperature values. We require that $\pi^{\beta_{0}}(\mathbf{x})=\pi(\mathbf{x})$ (that is, that $\beta_{0}=1$). Typically, the tempered target densities are simply powers of the original density, and we assume this for all parallel tempering experiments. Then, we run a Markov chain for each of the $n+1$ values of $\beta$, each with its own tempered target density. Eventually, the $\beta_{0}=1$ chain, also called the “cold” chain, with its stationary density $\pi^{\beta_{0}}(\mathbf{x}_{0})$ should correspond to the original target density of interest $\pi(\mathbf{x})$, and the idea is that using information from other chains with higher temperatures (lower $\beta$) will speed up the mixing in the cold chain.
@@
4.2 Optimal Scaling for Parallel Tempering
@@
    The target density is an i.i.d. product $\pi(\mathbf{x})=\prod_{i=1}^{d}f(x_{i})$ where $f$ is a single-dimensional component density. There are other smoothness and regularity restrictions on the target density taken from the main results of Roberts et al. [30], Roberts and Rosenthal [34].
@@ 5.3 Results
The empirical optimal swap acceptance rates for the three target distributions are summarized in Table 6, and sample ESJD curves for $d=30$ shown in Figure 8. The ESJD curves for PT are visibly more jagged than their RWM counterparts. This might be because the estimation of ESJD in this context is more susceptible to Monte Carlo error even with a large number of simulation steps, since the efficiency or acceptance rate of a swap move depends on the joint stationary distribution of two adjacent chains, and this is repeated for each temperature construction. Another possible reason might be the determinism in the way $\rho_{n}$ changes. Future work could investigate alternative algorithms to compute a fixed discrete temperature ladder that provides a certain swap rate prior to running the PT algorithm, perhaps by computing each change in $\rho_{n}$ from a distribution. Nonetheless, clear trends emerge from the results.
@@
Notably, the Rough Carpet target ($\pi_{4}$) shows a deviation. The optimal swap acceptance rate is consistently higher than 0.234 and increases away from the theoretical value with increasing dimension. This is surprising because the Rough Carpet target does satisfy the i.i.d. product form and other conditions of Assumption 4.1 by Atchadé et al. [1].
There are two plausible explanations that we favour here. The failure of the 0.234 rule here could be because the extreme multimodality of the target could reduce the practical utility of the ”hot” chains. The target possesses $3^{d}$ modes, and its probability mass is spread thinly across this vast landscape. For a hot chain where the inverse temperature $\beta$ is close to 0, the tempered density becomes extremely flat and loses the slight peak that a unimodal target would have. Consequently, it fails to efficiently sample from the specific modal regions that are of interest to the colder chains; the information flow from hot to cold chains is broken. To swap successfully more often, the temperature steps must be made smaller, which forces the swap acceptance rate to be higher.
An alternative and more simple explanation is that, due to the jaggedness of the ESJD vs swap acceptance rate trend in Fig 8, the discrete temperature ladder is only approximately optimal, and the true empirically optimal swap rate is somehow confounded by determinism in the ladder construction algorithm. As mentioned previously, alternative methods to construct temperature ladders yielding certain swap acceptance rates prior to running a PT algorithm could be an area for future work.
@@ 6 Discussion
We have presented an extensive empirical investigation into the generalizability of the 0.234 acceptance rate heuristic for both Random-Walk Metropolis (RWM) and Parallel Tempering (PT) algorithms, using a variety of target and proposal distributions. Our results demonstrate that the 0.234 rule’s utility extends beyond the classical i.i.d. product setting of Roberts et al. [30]. The heuristic remains approximately optimal for: (i) non-Gaussian Laplace and Uniform proposal distributions that satisfy a “shell” condition with a less stringent symmetry requirement than spherical symmetry; (ii) multimodal targets such as the “rough carpet” i.i.d product with $3^{d}$ modes, and non-product targets constructed as a sum of isotropic components, such as our Three-Mixture model; and (iii) even for highly correlated, non-linear targets like the Full and Hybrid Rosenbrock in sufficiently high dimensions ($d\geq 20$). These findings provide empirical validation for the generalized theory of Sherlock [39], which predicts that as long as the target’s geometry is sufficiently regular and isotropic in the high-dimensional limit, the 0.234 rule should hold. Our results in Section 3.2 even suggest that weaker conditions on the symmetry of the proposal than those required by Sherlock [39] could be sufficient.
@@
That being said, there are cases where 0.234 does not seem as relevant. The main case is in very low dimensions; in all our experiments with varying target densities, the optimal acceptance rate strays from 0.234 below a certain threshold which varies for each target density. In particular, targets composed of independent, low-dimensional anisotropic blocks (Even Rosenbrock) or those that are globally anisotropic in low dimensions (2D Rosenbrock), have an optimal acceptance rate significantly below 0.234. In these cases, we hypothesize the high-dimensional averaging effect that stabilizes the target’s geometry has not occurred, and the sampler’s performance is dictated by the challenging local geometry. As the dimension increases, the law of large numbers mitigates the effect of any single scaling factor or local correlation, so the target becomes effectively less eccentric, complying with the theory of Sherlock [39], and the optimal acceptance rate recovers towards 0.234. On the other hand, the continuous hypercube and Neal’s funnel completely fail the 0.234 heuristic. Although the hypercube is continuously differentiable everywhere except on its boundary, which has zero measure, an intuitive argument is that as $d$ increases, the probability that the proposal has all the values inside the hypercube decreases. As for the funnel, no single proposal scaling can be globally efficient, which is just a limitation of the RWM algorithm.
@@
Next, we discuss the optimal swap acceptance rate figure for parallel tempering and compare it with the optimal acceptance rate for random-walk Metropolis. On one hand, the results for the multivariate Gaussian and three-mixture density examples show that the 0.234 swap acceptance rate figure may be optimal even in lower dimensions. Although costly, if the practitioner is willing and able to construct the inverse temperature ladder iteratively with a spacing given by a swap acceptance probability, the spacing dictated by a 0.234 swap acceptance probability may be optimal for the practitioner. However, the theoretical 0.234 figure may not always be applicable: despite satisfying the i.i.d. conditions of the PT optimal scaling theory [1], the extreme multimodality ($3^{d}$ modes) of the “rough carpet” product density appears to render the hot chains ineffective in our experiments, with the caveat that the apparent failure of the 0.234 rule could partly be due to the jaggedness of our experimental ESJD vs swap acceptance rate curves. One hypothesis for this observed failure is that because the probability mass is spread so thinly, the hot chains fail to sample from meaningful modal regions, breaking the flow of information to the cold chain. This forces the algorithm to use smaller temperature steps (and thus higher swap acceptance rates) to achieve mixing, representing a practical failure of the 0.234 PT rule.
@@
Acknowledgements
Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute. We also thank the anonymous reviewer(s) for their helpful comments on earlier versions of this paper.